Sell or Process Further

Abica Coffee Company produces Columbian coffee in batches of 5,900 pounds. The standard quantity of materials required in the process is 5,900 pounds, which cost $6 per pound. Columbian coffee can be sold without further processing for $9 per pound. Columbian coffee can also be processed further to yield Decaf Columbian, which can be sold for $12 per pound. The processing into Decaf Columbian requires additional processing costs of $11,638 per batch. The additional processing also causes a 5% loss of product due to evaporation.

a. Prepare a differential analysis dated October 6 on whether to sell regular Columbian (Alternative 1) or process further into Decaf Columbian (Alternative 2). For those boxes in which you must enter subtracted or negative numbers use a minus sign.

Differential Analysis
Sell Regular Columbian (Alt. 1) or Process Further into Decaf Columbian (Alt. 2)
October 6
Sell
Regular
Columbian
(Alternative 1)
Process
Further into
Decaf
Columbian
(Alternative 2)
Differential
Effect
on Income
(Alternative 2)
Revenues $ $ $
Costs
Income (Loss) $ $ $

b. Should Abica sell Columbian coffee or process further and sell Decaf Columbian?

c. Determine the price of Decaf Columbian that would cause neither an advantage nor a disadvantage for processing further and selling Decaf Columbian. Round your answer to two decimal places.
$ per pound
Answer #1
Decaf Columbian units 5605 =5900*(1-5%)
a
Differential Analysis
Sell Regular Columbian (Alt. 1) or Process Further into Decaf Columbian (Alt. 2)
October 6
Sell Process Differential
Regular Further into Effect
Columbian Decaf on Income
(Alternative 1) Columbian (Alternative 2)
(Alternative 2)
Revenues 53100 67260 14160
Costs -35400 -47038 -11638
Income (Loss) 17700 20222 2522
b
Process further and sell Decaf Columbian
c
Price of Decaf Columbian 11.55 =12-(2522/5605)
Workings:
Sell Process
Regular Further into
Columbian Decaf
(Alternative 1) Columbian
(Alternative 2)
Revenues =5900*9 =5605*12
Costs =5900*6 =35400+11638